OpenStudy (saylilbaby):

what is a simplified trigonometric expression for csc 0 sin 0-sin^2 0? A.sec^20 B.cos^2 0 C.1-tan^2 0 D.tan0-sin^2 0

OpenStudy (danjs):

write everything in terms of sin and cos first

OpenStudy (danjs):

csc(a) = 1/sin(a)

OpenStudy (saylilbaby):

For the function y=tan 0, which angle measure has the same value as 2pie/3? A.5pie/3 B.7pie/6 C.4pie/3 D.pie/6

OpenStudy (danjs):

so you get 1 - sin^2(a)

OpenStudy (danjs):

and recall sin^2 + cos^2 = 1 ,

OpenStudy (danjs):

so, it is the same as cos^2(a)

OpenStudy (danjs):

remember the tangent has a period of pi

OpenStudy (saylilbaby):

what are the period and amplitude of the function y=2sin4 0? A.pie/2 and 2 B.pie and 4 C.pie /2 and 4 D.pie/2 and 1

OpenStudy (danjs):

add or subtract any multiple of pi

OpenStudy (saylilbaby):

so 1 is B @DanJS

OpenStudy (danjs):

yeah, just have to rewrite csc as 1/sin instead, then you get 1 - sin^2

OpenStudy (danjs):

which is an identity, sin^2 + cos^2 = 1

OpenStudy (saylilbaby):

is 2 B? @DanJS

OpenStudy (danjs):

tangent has a period of pi, rotation by pi will produce the same value

OpenStudy (saylilbaby):

so A @DanJS

OpenStudy (danjs):

2/3 pi + pi 2/3 pi + 2pi . . .

OpenStudy (saylilbaby):

c @DanJS

OpenStudy (danjs):

yeah A works there, 2/3 pi + pi = 5/3 pi

OpenStudy (danjs):

tangent of 2/3 pi = tangent of 5/3 pi try it

OpenStudy (saylilbaby):

i though c cause 2*2 is 4 and the denominator stays the same @DanJS

OpenStudy (danjs):

not doubling the angle, adding pi to it. 3/3 pi

OpenStudy (danjs):

2/3 pi + 3/3 pi = 5/3 pi

OpenStudy (danjs):

none of the others work, keep adding pi 8/3 pi 11/3 pi

OpenStudy (saylilbaby):

i see now okay next is 3 B @DanJS

OpenStudy (danjs):

y = 2*sin^4(a)

OpenStudy (danjs):

the amplitude is usually that number out front, A*sin(x), but here there is one difference

OpenStudy (saylilbaby):

by how much does the mean of the data set below change if the outlier is removed? round your answer to the nearest tenth, if necessary. 8,6,9,12,19,7,10,8 A.0.5 B.0.6 C.1.3 D.2.4

OpenStudy (danjs):

sin(x) is always between -1 and 1 (sin(x))^4 is always possitive because of the even power, so sin^4(x) is between 0 and 1

OpenStudy (saylilbaby):

so is it B. pie and 4 @DanJS

OpenStudy (danjs):

which one is the function \[y = \sin^4(x)\] or \[y = \sin(4x)\]

OpenStudy (danjs):

what are the period and amplitude of the function y=2sin4 0? A.pie/2 and 2 B.pie and 4 C.pie /2 and 4 D.pie/2 and 1

OpenStudy (saylilbaby):

a @DanJS

OpenStudy (danjs):

is the 4 a power or part of the angle

OpenStudy (danjs):

the first power one, amplitude is 1 period is pi second one is amplitude 2, period T=2pi/4 = pi/2

OpenStudy (danjs):

here is a little thing on the general form for trig function http://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.754992.html

OpenStudy (saylilbaby):

so its A @DanJS

OpenStudy (danjs):

but the 4 power keeps the sin function positive, so th eamplitude is not 2, it would be half of it, only positive values

OpenStudy (danjs):

if the function is y = 2*sin(4x) , not sin to the 4th power, then the answer is A yeah

OpenStudy (saylilbaby):

c.pie/2 and4 @DanJS

OpenStudy (danjs):

if you have y = A*sin(Bx) the amplitude is A the period is (2pi)/B

OpenStudy (danjs):

for y = 2*sin(4x) period is (2pi)/4 or pi/2 amplitude is 2

OpenStudy (saylilbaby):

a,b,c or d @DanJS